A cylinder is within the cube touching all the vertical face. A cone is inside the cylinder. If their height are the same with the same base find the ratio of their volume.
 

Grade:9

1 Answers

Arun
4 years ago
Assume that the length of all edges of the cube=x units
Formula for volume of cube=x³ cube unit
Due an occurrence that the cylinder is within the cube &it touches all the vertices faces of the cube.
Therefore radius of the base of the cylinder & cone & height of cylinder of a cone=x
Volume of the cylinder=π×{x/2}²×x=22/7×x³/4
11/14x³ cube unit
Volume of cone should be=1/3π{x/2)²×x=11/42 cubic unit
Required ratio=volume of cube:volume of cylinder:volume of cone
=x³:11/14x³:11/42x³=42:33:11